What is the slope for these problems, I forgot how to do these.
1. 5-2(x+5)>30
2. 3x-1/2(6x+4)=6-2x
3. 12x+5=6x=17
4. -x/5+12>14

Answer by Boreal(15235)

5-2(x+5)>30
5-2x-10>30
-5-2x>30
-2x>35
x<-17.5, change inequality direction when dividing by a - number.
2. 3x-1/2(6x+4)=6-2x
3x-3x-2=6-2x
2x-2=6
2x=8
x=4
12-(1/2)(28)=6-8
12-14=-2

3. 12x+5=6x+17; I edited a "+"
6x=12
x=2
24+5=12+17 check
4. -x/5+12>14
-x/5>2
multiply everything by -5 and change inequality
x<-10
check with -15
3+12>14 check.